### v1

@@ Abstract
The Axelrod approximation is widely used in astrophysical modelling codes to evaluate electron-impact excitation effective collision strengths for forbidden transitions. Approximate methods such as this are a necessity for many heavy elements with open shells where collisional data is either non existent or sparse as the use of more robust methods prove prohibitively expensive. Atomic data for such forbidden transitions are essential for producing full collisional radiative models that do not assume Local-Thermodynamic-Equilibrium (LTE). In this short work we repeat the optimization of the simple Axelrod formula for a large number of -matrix data sets, ranging from Fe and Ni to the first r-process peak elements of Sr, Y and Zr, to higher Z systems Te, W, Pt and Au. We show that the approximate treatment of forbidden transitions can be a significant source of inaccuracy in such collisional radiative models. We find a large variance of the optimized coefficients for differing systems and charge states, although some general trends can be seen based on the orbital structure of the ground-state-configurations. These trends could potentially inform better estimates for future calculations for elements where -matrix data is not available.
@@ 1 Introduction
The accurate calculation of atomic data for collision processes is computationally expensive and typically has to be carried out on a case-by-case basis for each ion and ionization stage of interest. For this reason, in models where the approximation of Local-Thermodynamic-Equilibrium (LTE) cannot be assumed, the employed codes typically make use of approximate or semi-empirical datasets. The formulae of Axelrod [1] and van Regemorter [2] are often used when calculating the rates of electron-impact-excitation. Examples of works making use of these approximations include [3, 4, 5]. While such formulae are valid for producing usable data for the simulation of astrophysical transients, it has been shown recently by Bromley et al. [7] and Mulholland et al. [8] that the data these methods produce in their current state are typically inaccurate for many transitions. Generally for forbidden lines the collision strengths are underestimated by several orders of magnitude when compared with results from the -matrix method of Burke [6]. Additionally, at low temperatures the -matrix method is more reliable due to the inclusion of resonances in the low energy interaction region. The inclusion of resonances is dependent on the specific atomic system, which increases computational complexity - however their neglection can potentially reduce the calculated rates by a significant margin.
@@
Recently, a large amount of atomic data has become available for heavy elements of astrophysical interest [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 7], particularly those for kilonovae (KNe), supernovae (SNe) and fusion applications. In this work we compare optically thin emission spectra calculated using the Axelrod [1] and van Regemorter [2] empirical formulae with models based on data produced by sophisticated -matrix calculations. In particular, the validity of the Axelrod formula for forbidden transitions in the context of modelling KNe is investigated. Use of the Axelrod and van Regemorter formulae has been discussed previously [7, 22, 8], and some revision has also been put forward by Pognan et al. [4]. In this publication we add to these discussions by presenting an ion-dependent re-optimization of the Axelrod formula based on the availability of the new -matrix data.
@@
In Section 2, the theory of collisional-radiative-modelling and electron-impact-excitation is briefly introduced, as well as the approximate formulae of van Regemorter [2] and Axelrod [1]. It is also demonstrated, by way of an illustrative example, how the Axlerod formula in its unaltered form can systematically under predict emission from ions. In Section 3, we detail the re-optimization of the coefficients in the Axelrod formula, where we optimize toward newly available calibrated -matrix data. Optically thin emission spectra are constructed using a selection of datasets to show how the Axelrod formula can be amended by multiplying the coefficients by an ion-dependent scaling factor. Some basic trends in terms of the shell-behaviour of the ions considered are further analysed. In Section 4 we conclude with a summary of the main findings of the paper and an outlook.
@@ 2 Theory
where the populations are appropriately normalised, is the electron density, and is the Photon-Emissivity-Coefficient for the transition. More details surrounding collisional radiative modelling can be found in Summers et al. [23], Burke [6].
@@
where , is an effective Gaunt factor tabulated in [2] for neutral and singly ionized species, is the relevant Einstein A-coefficient, is a statistical weight, and is the wavelength of the transition in units of cm. This formula was given in van Regemorter [2] for neutral and singly ionized species. For species more ionized than this one can make use of other approximations such as those detailed in the ADAS routine d7pyvr [24]. From [samspon92] (they use as lower weight)
Using ( is upper)
Gives:
Put and integrate both sides to arrive at eq. 6. So I conclude it should probably be the upper weight.
@@ 3.2 Results
The Axelrod approximation was optimised for a number of atomic species for which R-matrix data was readily available. Reoptimized coefficients are shown in Table 1. The values of are those that minimized the function in Eq. (9) and the values of are those that minimized the function in Eq. (10). Elements are grouped by the highest orbital angular momentum in their ground state configuration222Except for Y iii, which is grouped with the s-shell elements due to its effective-single-electron structure.. It can be seen that quite generally is larger than for each element, in some cases by an order of magnitude. Given that the Axelrod formula is typically used for the ad hoc representation of the general behaviour within a collisional radiative model, the larger values found in will more accurately represent the cooling. It is seen from Table 1 that the -shell elements generally require a factor 10 - 100 times larger than that of the standard Axelrod formula. Similarly, -shell elements require a factor around 10-20 times higher. By contrast, the -shell elements appear well represented in order of magnitude by the stock formula. This is perhaps expected given the Axelrod formula [1] was originally optimized on the Fe iii (3d6) data of Garstang et al. [25]. With expanding atomic data sets, it will be interesting to see the impact of -shell elements when large amounts of atomic data become available for them.
@@
-shell | -shell | cont | |||||||
Sr i | 4p6 5s2 | 0.05366 | 0.02609 | [9] | Ni ii | 3d9 | 0.00290 | 0.00037 | [14] |
Y ii | 4p6 5s2 | 0.04460 | 0.03247 | [8] | Ni iii | 3d8 | 0.00668 | 0.00201 | [15] |
Sr ii | 4p6 5s | 0.47730 | 0.21053 | [8] | Ni iv | 3d7 | 0.00472 | 0.00121 | [16] |
Y iii | 4p6 4d | 0.37418 | 0.23215 | [11] | Zr i | 4d2 5s2 | 0.00503 | 0.00212 | [17] |
Zr ii | 4d2 5s | 0.01674 | 0.00631 | [17] | |||||
-shell | Zr iii | 4d2 | 0.05006 | 0.02053 | [17] | ||||
Sr iii | 4p6 | 0.07666 | 0.04206 | [9] | W i | 6s2 5d4 | 0.00283 | 0.00139 | [18] |
Te i | 5s2 5p4 | 0.06736 | 0.01168 | [10] | W ii | 6s2 5d3 | 0.00372 | 0.00203 | [19] |
Te ii | 5s2 5p3 | 0.07460 | 0.05840 | [10] | W iii | 6s2 5d2 | 0.00653 | 0.00431 | [20] |
Te iii | 5s2 5p3 | 0.08868 | 0.06478 | [10] | Pt i | 5d9 6s | 0.00543 | 0.00148 | [7] |
Pt ii | 5d9 | 0.00444 | 0.00174 | [7] | |||||
-shell | Pt iii | 5d8 | 0.00309 | 0.00147 | [7] | ||||
Fe ii | 3d6 4s | 0.00233 | 0.00058 | [12] | Au i | 5d10 6s | 0.01183 | 0.00554 | [21] |
Fe iii | 3d6 | 0.00274 | 0.00075 | [13] | Au ii | 5d10 | 0.00471 | 0.00055 | [21] |
Au iii | 5d9 | 0.00373 | 0.00176 | [21] |
@@
We additionally show on Figure 3 the scaled emission for Au i, employing a similar set of modified data to the above. Au i presents a special case of a -shell element whose electronic configurations are constructed from the 5d10 and 5d9 series. The former set behaves like a single electron system. There is therefore a limited number of forbidden transitions that are typically stronger than those of other systems. For this reason the optimized Axelrod coefficient is considerably higher than the surrounding -shell elements presented in Table 1. Figure 3 therefore demonstrates a mixed bag of agreement in the emission estimations because the Axelrod formula assumes that all the forbidden lines behave the same, which is evidently not the case for this ion.
@@
By contrast (and is the case for the majority of the lines shown) there is significant underestimation of the emission in the case of the base Axelrod approximation, where agreement is seen between the data sets employing this with either the R-matrix or van Regemorter allowed rates. Moving to the re-optimized Axelrod approximation introduces a certain improvement in the estimation of the emission, in many cases by an order of magnitude. The main victory is the improvement of the evidently strongest lines likely to contribute to a real astrophysical spectrum.
@@ 4 Conclusions and Outlook
To summarise, the Axelrod approximation for effective collision strengths has been reviewed and revisited. The Axelrod coefficient has been calculated for a variety of ion stages with available R-matrix data. The elements studied cover the 1st (Sr-Zr), 2nd (Te) and 3rd (W,Pt,Au) peaks of the rapid nucleosynthesis processes. The values presented here could in principle be used to generate collision strengths for collisions between high-lying levels not included in structure calculations, or aid in the calculation of similarly structured elements by using ball-park coefficients reflective of those presented in Table 1. Elements with -shell ground states tend to have transitions 10 - 1000 times stronger than the base Axelrod approximation. -shell ground states tend to be 10-20 times stronger. -shell elements tend to have forbidden transitions of the rough order of magnitude of the original Axelrod approximation. Despite this, the relatively large variance in indicates that the Axelrod approximation should be reserved for where no data is available, and should be replaced with -matrix data as it comes available. Nonetheless, it is intended that this work serves as an intermediate step in the improvement of atomic data for astrophysical models, where the values presented here can be used to provide more information for elements with no data available and the Axelrod formula can be scaled based on the electronic configurations of the species involved.

### v2

@@ Abstract
The Axelrod approximation is widely used in astrophysical modelling codes to evaluate electron-impact excitation effective collision strengths for forbidden transitions. Approximate methods such as this are a necessity for many heavy elements with open shells where collisional data is either non existent or sparse as the use of more robust methods prove prohibitively expensive. Atomic data for such forbidden transitions are essential for producing full collisional radiative models that do not assume Local-Thermodynamic-Equilibrium (LTE). In this short work we present the re-optimization the simple Axelrod formula for a large number of -matrix data sets, ranging from Fe and Ni to the first r-process peak elements of Sr, Y and Zr, to higher systems Te, W, Pt and Au. We show that the approximate treatment of forbidden transitions can be a significant source of inaccuracy in such collisional radiative models. We find a large variance of the optimized coefficients for differing systems and charge states, although some general trends can be seen based on the orbital structure of the ground-state-configurations. These trends could potentially inform better estimates for future calculations for elements where -matrix data is not available.
@@ 1 Introduction
The accurate calculation of atomic data for collision processes is computationally expensive and typically has to be carried out on a case-by-case basis for each ion and ionization stage of interest. For this reason, in models where the approximation of Local-Thermodynamic-Equilibrium (LTE) cannot be assumed, the employed codes typically make use of approximate or semi-empirical datasets. The formulae of Axelrod [1] and van Regemorter [2] are often used when calculating the rates of electron-impact-excitation. Examples of works making use of these approximations include [3, 4, 5]. While such formulae are valid for producing usable data for the simulation of astrophysical transients, it has been shown recently by Bromley et al. [7] and Mulholland et al. [8] that the data these methods produce in their current state are typically inaccurate for many transitions. Generally for forbidden lines the collision strengths are underestimated by several orders of magnitude when compared with results from the -matrix method of Burke [6]. Additionally, at low temperatures the -matrix method is more reliable when compared with methods such as distorted-wave due to the inclusion of resonances in the low energy interaction region. The inclusion of resonances is dependent on the specific atomic system, which increases computational complexity - however their neglection can potentially reduce the calculated rates by a significant margin.
@@
Recently, a large amount of atomic data has become available for heavy elements of astrophysical interest [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 7], particularly those for kilonovae (KNe), supernovae (SNe) and fusion applications. In this work we compare optically thin emission spectra calculated using the Axelrod [1] and van Regemorter [2] empirical formulae with models based on data produced by sophisticated -matrix calculations. In particular, the validity of the Axelrod formula for forbidden transitions in the context of modelling KNe is investigated. Use of the Axelrod and van Regemorter formulae has been discussed previously [7, 22, 8], and some revision has also been put forward by Pognan et al. [4]. In this publication we add to these discussions by presenting an ion-dependent re-optimization of the numerical constant within the Axelrod formula based on the availability of the new -matrix data. The work here serves to illustrate the degree in variation in characteristic collision strengths between different ions, and quantify the extent to which results from full calculations can be captured with such simple formulae. We also discuss how existing data can serve to guide better informed approximation for elements where no -matrix data is available.
@@
In Section 3, the theory of collisional-radiative-modelling and electron-impact-excitation is briefly introduced, as well as the approximate formulae of van Regemorter [2] and Axelrod [1]. It is also demonstrated, by way of illustrative examples, how the Axelrod formula in its unaltered form can systematically under predict emission from ions. In Section 4, we detail the re-optimization of the coefficients in the Axelrod formula, where we optimize toward newly available calibrated -matrix data. Optically thin emission spectra are constructed using a selection of datasets to show how the Axelrod formula can be amended with an ion-dependent coefficient. Some basic trends in terms of the shell-behaviour of the ions considered are further analysed. In Section 5 we conclude with a summary of the main findings of the paper and an outlook.
2 Theory
Collisional radiative modelling codes solve the rate equations for the level populations of ions in a plasma,
where the matrix encapsulates the rates of the atomic processes involved. This includes the electron-impact excitation and de-excitation rates given by
where is the electron temperature in Kelvin, is the Boltzmann constant, is the statistical weight of the lower (upper) level and is the energy difference between the levels. The quantities represent Maxwellian-averaged-collision strengths. These have been calculated a number of ways, including distorted wave [23] and the R-matrix codes [6]. As discussed in the introduction, in lieu of data produced by these methods, semi-empirical formulae are employed, such as those presented in [2, 1]. The focus of this work is on forbidden lines, for which modelling codes typically employ the formula of [1],
which was optimized on data from [24]. This coefficient of 0.004 was specifically produced for data of Fe, and will be shown to result in a systematic underestimation of atomic line emission. The presented formula is independent of temperature, which is discussed in [25].
@@ 3 Theory
where the populations are appropriately normalised, is the electron density, and is the Photon-Emissivity-Coefficient for the transition. More details surrounding collisional radiative modelling can be found in Summers2006, Burke [6].
@@
where , is an effective Gaunt factor tabulated in [2] for neutral and singly ionized species, is the relevant Einstein A-coefficient, is a statistical weight, and is the wavelength of the transition in units of cm. This formula was given in van Regemorter [2] for neutral and singly ionized species. For species more ionized than this one can make use of other approximations such as those detailed in the ADAS routine d7pyvr [openadas_manual].
@@ 4.2 Results
The Axelrod approximation was optimised for a number of atomic species for which R-matrix data was readily available. Reoptimized coefficients are shown in Table 1. The values of are those that minimized the function in Eq. (13) and the values of are those that minimized the function in Eq. (14). Elements are grouped by the highest orbital angular momentum in their ground state configuration. It can be seen that quite generally is larger than for each element, in some cases by an order of magnitude. Given that the Axelrod formula is typically used for the ad hoc representation of the general behaviour within a collisional radiative model, the larger values found in will more accurately represent the cooling. It is seen from Table 1 that the -shell elements generally require a factor 10 - 100 times larger than that of the standard Axelrod formula. Similarly, -shell elements require a factor around 10-20 times higher. By contrast, the -shell elements appear well represented in order of magnitude by the stock formula. This is perhaps expected given the Axelrod formula [1] was originally optimized on the Fe iii (3d6) data of Garstang et al. [24]. With expanding atomic data sets, it will be interesting to see the impact of -shell elements when large amounts of atomic data become available for them. Additionally, we have presented the mean values of within the table, for a more universal replacement in models for ions where accurate data is not yet available. Use of these mean values should come with caution, as there remains large spread within the shell groups.
@@
| -shell | 2.4 | -shell | continued | ||||||
| Sr i | 4p6 5s2 | 5.4 | 2.6 | [9] | Ni ii | 3d9 | 2.9 | [14] | |
| Y ii | 4p6 5s2 | 4.5 | 3.2 | [8] | Ni iii | 3d | 6.7 | [15] | |
| Sr ii | 4p6 5s | 4.8 | 2.1 | [8] | Ni iv | 3d | 4.7 | [16] | |
| Y iii | 4p6 4d | 3.7 | 2.3 | [11] | Zr i | 4d 5s2 | 5.0 | [17] | |
| Zr ii | 4d2 5s | 1.7 | [17] | ||||||
| -shell | 7.7 | Zr iii | 4d2 | 5.0 | [17] | ||||
| Sr iii | 4p6 | 7.7 | 4.2 | [9] | W i | 6s2 5d4 | 2.8 | [18] | |
| Te i | 5s2 5p4 | 6.7 | 1.2 | [10] | W ii | 6s2 5d3 | 3.7 | [19] | |
| Te ii | 5s2 5p3 | 7.5 | 5.8 | [10] | W iii | 6s2 5d2 | 6.5 | [20] | |
| Te iii | 5s2 5p2 | 8.9 | 6.5 | [10] | Pt i | 5d9 6s | 5.4 | [7] | |
| Pt ii | 5d9 | 4.4 | [7] | ||||||
| -shell | 8.1 | Pt iii | 5d8 | 3.1 | [7] | ||||
| Fe ii | 3d6 4s | 2.3 | 5.8 | [12] | Au i | 5d10 6s | 1.2 | [21] | |
| Fe iii | 3d6 | 2.7 | 7.5 | [13] | Au ii | 5d10 | 4.7 | [21] | |
| Au iii | 5d9 | 3.7 | [21] |
@@
We additionally show on Figure 3 the scaled emission for Au i, employing a similar set of modified data to the above. Au i presents a special case of a -shell element whose electronic configurations are constructed from the 5d10 and 5d9 series. The former set behaves like a single electron system. There is therefore a limited number of forbidden transitions that are typically stronger than those of other systems. For this reason the optimized Axelrod coefficient is considerably higher than the surrounding -shell elements presented in Table 1. Figure 3 therefore demonstrates a lack of consistency in terms agreement in the emission estimations because the Axelrod formula assumes that all the forbidden lines behave the same, which is evidently not the case for this ion.
@@
By contrast (and is the case for the majority of the lines shown) there is significant underestimation of the emission in the case of the base Axelrod approximation, where agreement is seen between the data sets employing this with either the R-matrix or van Regemorter allowed rates. Moving to the re-optimized Axelrod approximation introduces a certain improvement in the estimation of the emission, in many cases by an order of magnitude. The primary result is the improvement of the evidently strongest lines which are most likely to contribute to a real astrophysical spectrum.
@@ 5 Conclusions and Outlook
To summarise, the Axelrod approximation for effective collision strengths has been reviewed and revisited. The Axelrod coefficient has been calculated for a variety of ion stages with available R-matrix data. The elements studied cover the 1st (Sr-Zr), 2nd (Te) and 3rd (W,Pt,Au) peaks of the rapid nucleosynthesis processes. The values presented here could in principle be used to generate collision strengths for collisions between high-lying levels not included in structure calculations, or aid in the calculation of similarly structured elements by using approximate coefficients reflective of those presented in Table 1. Elements with -shell ground states tend to have transitions 10 - 1000 times stronger than the base Axelrod approximation. -shell ground states tend to be 10-20 times stronger. -shell elements tend to have forbidden transitions of the rough order of magnitude of the original Axelrod approximation. Despite this, the relatively large variance in indicates that the Axelrod approximation should be reserved for where no data is available, and should be replaced with -matrix data as it comes available. Nonetheless, it is intended that this work serves as an intermediate step in the improvement of atomic data for astrophysical models, where the values (or perhaps their mean values) presented here can be used to provide more information for elements with no data available and the Axelrod formula can be scaled based on the electronic configurations of the species involved.